At the beginning of the month, Arthur's Olde Consulting Corporation had two jobs in process that had the following costs assigned from previous months:

Job Number Direct Labor Applied Overhead
SY-400 $ 23,940 ?
SY-403 16,020 ?

During the month, Jobs SY-400 and SY-403 were completed but not billed to customers. The completion costs for SY-400 required $27,000 in direct labor. For SY-403, $81,000 in labor was used.

During the month, the only new job, SY-404, was started but not finished. Total direct labor costs for all jobs amounted to $151,020 for the month. Overhead in this company refers to the cost of work that is not directly traced to particular jobs, including copying, printing, and travel costs to meet with clients. Overhead is applied at a rate of 60 percent of direct labor costs for this and previous periods. Actual overhead for the month was $93,300.

Required:

a. What are the costs of Jobs SY-400 and SY-403 at the beginning of the month and when completed?

b. What is the cost of Job SY-404 at the end of the month?

c. How much was under- or overapplied service overhead for the month?
Answer #1

a.

Cost at the beginning:

Job SY-400 = $23,940 + ($23,940 X 60%) = $38,304

Job SY-403 = $16,020 + ($16,020 X 60%) = $25,632

Cost at the end:

Job SY-400 = $38,304 + [$27,000 + ($27,000 X 60%)] = $81,504

Job SY-403 = $25,632 + [$81,000 + ($81,000 X 60%)] =$155,232

b.

Cost of Job SY-404 = ($151,020 - $27,000 - $81,000) + 60% = $68,832

c.

Total overhead applied = $151,020 X 60% = $90,612

Actual overhead incurred is more then applied overhead, then overhead is underapplied.

= $93,300 - $90,612

= $2,688 underapplied